OpenStudy (anonymous):

change f(x) = 30(0.94)^x to an exponential function with base e.

OpenStudy (anonymous):

General form of exponential function is \[y= A e^{kx}\]

OpenStudy (anonymous):

Which, by index laws can be expressed as

OpenStudy (anonymous):

\[y=A[e^{k}]^{x}\]

OpenStudy (anonymous):

remember when you raise one power to another power, you multiply the powers.

OpenStudy (anonymous):

Now it's just a matter of matching terms.

OpenStudy (anonymous):

A=30 (obviously)

OpenStudy (anonymous):

e^(k) = 0.94 ---> k = ln(0.94) = -0.06188 (to 5dp)

OpenStudy (anonymous):

So, \[y= 30e^{-0.06188x}\]

OpenStudy (anonymous):

thank you so much! makes perfect sense